find the whole number such that four times the number subtracted from 3 times the square of the number makes 15.
Answer by mananth(16949)

let the number be x
3 times square = 3x^2
4 times the number = 4x
..
3x^2-4x=15
3x^2-4x-15=0
3x^2-9x+5x-15=0
3x(x-3)+5(x-3)=0
(x-3)(3x+5)=0
x=3 OR x=-5/3
x= 3 the whole number